# Reinforcing Infrastructure Networks with Multicriteria Portfolio Decision Analysis: An Application to Railway Stations in Finland

### v1

@@ Appendix A Upper Bound of the Portfolio Performance
An alternative approach to generate tighter bounds at a higher computational cost is presented in the following. The solution of the optimization problem (7) provides an upper bound $\hat{u_{w_{e}}}$ of the expected utility at the extreme $w_{e}$, achievable by implementing a feasible portfolio by adding more reinforcement actions to portfolio $q$. Where $C$ and $D$ are the indices of the actions implemented and not implemented in the portfolio $q$, i.e., $C=\{l\mid q_{l}=1\}$ and $D=\{l\mid q_{l}=0\}$.
@@
Let $q^{*}$ denote a portfolio whose implementation provides the upper bounds of the expected utility at each extreme point, computed by solving (7). If any cost-efficient portfolio dominates portfolio $q^{*}$, the portfolio $q$ can be discarded because no cost-efficient portfolio can be generated from it. The main advantage of this approach is that the bounds are computed by considering the remaining available resources. One can also incorporate logical constraints into the optimization problem. Nevertheless, solving the problem (7) can lead to a different portfolio for each extreme point. If that is the case, it means that $q^{*}$ is not necessarily a feasible portfolio, and consequently, the upper bounds are not necessarily exact. Since $q^{a}$ implies implementing all the remaining reinforcement actions, the upper bounds provided by $q^{*}$ are always lower or equal (worst case) than the ones provided by $q^{a}$.
The efficiency of the methods depends on several parameters, such as the utility function, available budget, cost of the reinforcement actions, etc. For example, optimization problems with complex utility functions may be hard to solve, making the computation of $\hat{u_{w_{e}}}$ impractical. In other cases, where the budget is small and just a few actions can be implemented, the bounds provided by $q_{a}$ can be very loose. One could also complement the methodologies by applying Algorithm 1 as presented and switch to the optimization model on the points where the portfolios $q^{a}$ are infeasible.

### v2

@@ Appendix A Upper Bound of the Portfolio Performance
An alternative approach to generate tighter bounds at a higher computational cost is presented. The solution of the optimization problem (7) provides an upper bound $\hat{u_{w_{e}}}$ of the expected utility at the extreme $w_{e}$, achievable by implementing a feasible portfolio by adding more fortification actions to portfolio $q$. Where $C$ and $D$ are the indices of the actions implemented and not implemented in the portfolio $q$, i.e., $C=\{l\mid q_{l}=1\}$ and $D=\{l\mid q_{l}=0\}$.
@@
Let $q^{*}$ denote a portfolio whose implementation provides the upper bounds of the expected utility at each extreme point, computed by solving (7). If any cost-efficient portfolio dominates portfolio $q^{*}$, the portfolio $q$ can be discarded because no cost-efficient portfolio can be generated from it. The main advantage of this approach is that the bounds are computed by considering the remaining available resources. One can also incorporate logical constraints into the optimization problem. Nevertheless, solving the problem (7) can lead to a different portfolio for each extreme point. If that is the case, it means that $q^{*}$ is not necessarily a feasible portfolio, and consequently, the upper bounds are not necessarily exact. Since $q^{a}$ implies implementing all the remaining fortification actions, the upper bounds provided by $q^{*}$ are always lower or equal (worst case) than the ones provided by $q^{a}$.
The efficiency of the methods depends on several parameters, such as the utility function, available budget, cost of the fortification actions, etc. For example, optimization problems with complex utility functions may be hard to solve, making the computation of $\hat{u_{w_{e}}}$ impractical. In other cases, where the budget is small and just a few actions can be implemented, the bounds provided by $q_{a}$ can be very loose. One could also complement the methodologies by applying Algorithm 1 as presented and switch to the optimization model on the points where the portfolios $q^{a}$ are infeasible.